### v1

# Bethe equations for the critical three-state Potts spin chain with toroidal boundary conditions

Bethe equations for the critical three-state Potts spin chain with toroidal boundary conditions
M.J. Martins
Universidade Federal de São Carlos
Departamento de Física
C.P. 676, 13565-905, São Carlos (SP), Brazil
In this paper, we consider the parameterization of the spectra of the three-state critical Potts quantum chain with integrable twisted boundary conditions in terms of Bethe ansatz type equations. The Bethe equations are found by investigating the structure of the eigenvalues of the respective twisted transfer matrices, and with the help of certain identities satisfied by the product of transfer matrices operators. We have studied the completeness of the spectrum in terms of the Bethe roots for small lattice sizes and have computed the eigenstate momenta. We found that the spins of the low-lying excitations can have fractional values in accordance with predictions of the underlying conformal field theory. We argue that our framework can be used to build integrable Hamiltonians whose spectra are determined by mixing different toroidal boundary conditions.
Keywords: three-state Potts quantum chain, twisted boundaries, Bethe equations.
December  2025
1 Introduction
In general, integrable classical lattice spin models of statistical mechanics are usually considered by imposing periodic boundary conditions on the two-dimensional lattice [1]. However, it is expected that integrability will remain intact for the other boundary conditions compatible with the torus and the system should be solvable under more generic twisted boundary conditions. The typical example is the $Z(2)$ invariant Ising model which is considered to be exactly solvable for both periodic and antiperiodic boundary conditions [2]. In this case, the antiperiodic boundary is equivalent of changing the sign of the exchange interaction at one end of the toroidal lattice. The purpose of this paper is to consider this problem for one of the simplest generalization of the Ising model, where the spin variables can assume three possible values. This is the so-called three-state scalar Potts model in which the interactions among adjacent spins are assumed to have the same thermal energy [3]. On the square lattice, this system has only two horizontal and vertical edge weights, and it becomes solvable when such weights sit on a curve in which the model is self-dual [4]. We recall that such a manifold can be regarded as a special case of the $n$-state integrable $Z(n)$ invariant spin model discovered by Fateev and Zamolodchikov [5]. It turns out that the horizontal $W_{h}(x)$ and vertical $W_{v}(x)$ edge weights of the integrable three-state scalar Potts model have the following structure [5],
$W_{h}(x)=\left(\begin{array}[]{ccc}1&a(x)&a(x)\\ a(x)&1&a(x)\\ a(x)&a(x)&1\\ \end{array}\right),~~~W_{v}(x)=\left(\begin{array}[]{ccc}1&b(x)&b(x)\\ b(x)&1&b(x)\\ b(x)&b(x)&1\\ \end{array}\right),$ (1)
where the variable $x$ parameterizes the self-dual curve. The non-trivial matrix elements $a(x)$ and $b(x)$ are determined by,
$a(x)=\frac{\sin(\frac{\pi}{6}-x)}{\sin(\frac{\pi}{6}+x)},~~~b(x)=\frac{\sin(x)}{\sin(\frac{\pi}{3}-x)}.$ (2)
One consequence of integrability is that we can construct a family of commuting transfer matrices, and from that we derive the underlying quantum spin chain [1]. In the case of classical spin models, the suitable transfer matrix is built by combining the horizontal and vertical edge weights along the diagonals of a square $L\times L$ lattice following the original construction devised by Onsager for the Ising model [6]. This operator is usually denominated the diagonal-to-diagonal transfer matrix and its matrix elements are given by
$\left[T_{\mathrm{dia}}(x)\right]^{b_{1},\dots,b_{L}}_{a_{1},\dots,a_{L}}=\prod_{j=1}^{L}W_{v}(a_{j},b_{j}|x)W_{h}(a_{j},b_{j+1}|x),$ (3)
where $b_{L+1}=b_{1}$ for periodic boundary conditions. For the three-state scalar Potts model, the elements $W_{h,v}(i,j|x)$ are the entries of the matrices defined by Eqs.(1,2).
The respective three-state quantum spin chain is obtained by expanding the diagonal-to-diagonal transfer matrix (3) around the point $x=0$, see for instance [7, 8, 9]. At this point the transfer matrix reduces to the identity matrix and the next term in the expansion gives rise to the Hamiltonian operator. On a chain of $L$ sites with periodic boundary conditions, the Hamiltonian can be written as
$H^{(p)}=-\frac{2}{\sqrt{3}}\sum_{j=1}^{L-1}\left(Z_{j}Z^{\dagger}_{j+1}+Z^{\dagger}_{j}Z_{j+1}+X_{j}+X^{\dagger}_{j}\right)-\frac{2}{\sqrt{3}}\left(Z_{L}Z^{\dagger}_{1}+Z^{\dagger}_{L}Z_{1}+X_{L}+X^{\dagger}_{L}\right),$ (4)
where the operators $Z_{j}$ and $X_{j}$ acting at the $j$-th site of the chain obey the $Z(3)$ algebra. Their explicit matrix representations are,
$Z=\left(\begin{array}[]{ccc}1&0&0\\ 0&\omega&0\\ 0&0&\omega^{2}\\ \end{array}\right),~~X=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\\ \end{array}\right),~~~\omega=\exp(2\pi i/3).$ (5)
The typical feature of an integrable quantum spin chain is that its spectrum can be parametrized in terms of a set of nonlinear relations usually denominated Bethe ansatz equations [10]. This problem for the Hamiltonian (4) was studied by Albertini [11] who has derived a Bethe equation for the respective eigenvalues, exploring certain matrix identities for products of diagonal-to-diagonal transfer matrices [12, 13]. For comparison with our results for other toroidal boundary conditions, we summarize in the following the main conclusions of the work [11]. The Hamiltonian (4) is invariant by the $Z(3)$ and its Hilbert space can be separated into three sectors labeled by the $Z(3)$ charge $\exp(\frac{i\pi Q}{3})$ where $Q=0,1,2$. The respective Bethe equation for a given sector is [11]
$\bigg[\frac{\sinh(\lambda_{j}+i\frac{\pi}{12})}{\sinh(\lambda_{j}-i\frac{\pi}{12})}\bigg]^{2L}=(-1)^{L}\prod_{\stackrel{{\scriptstyle k=1}}{{k\neq j}}}^{N_{Q}}\frac{\sinh(\lambda_{j}-\lambda_{k}+i\frac{\pi}{3})}{\sinh(\lambda_{j}-\lambda_{k}-i\frac{\pi}{3})},~~j=1,\dots,N_{Q},$ (6)
where the number of Bethe roots $N_{Q}$ for each sector is $N_{0}=2L$ and $N_{1}=N_{2}=2L-2$. The energies $E^{(p)}_{Q}$ of the Hamiltonian (4) are expressed in terms of the Bethe roots by,
$E^{(p)}_{Q}=\sum_{j=1}^{N_{Q}}\cot\Big(\frac{\pi}{12}-i\lambda_{j}\Big)-\frac{2L}{\sqrt{3}},$ (7)
and several consequences of such a Bethe ansatz solution were studied subsequently in [14, 15].
Before proceeding, we recall that the Hamiltonian (4) is in fact invariant under the larger permutation group $S_{3}$. This means that this operator commutes with the charges ${\cal{O}}_{Z(3)}$ and ${\cal{O}}_{Z(2)}$ associated with the implementation of the underlying symmetries $Z(3)$ and $Z(2)$, respectively. These two charges are defined by,
${\cal{O}}_{Z(3)}=\prod_{j=1}^{L}X_{j},~~{\cal{O}}_{Z(2)}=\prod_{j=1}^{L}C_{j},$ (8)
where the operator $C$ corresponds to complex conjugation and its matrix representation is,
$C=\left(\begin{array}[]{ccc}1&0&0\\ 0&0&1\\ 0&1&0\\ \end{array}\right).$ (9)
It turns out that the $S_{3}$ invariance can be broken by imposing suitable toroidal boundary conditions while preserving the integrability of the corresponding quantum spin chain. This feature, for instance, can be inferred considering the results for the operator content of a conformally invariant theory with the central charge realized by the critical three-state Potts quantum chain. This model is known to have $c=\frac{4}{5}$ [16, 17] and it has been argued that this class of universality is compatible with two other toroidal boundary conditions in addition to the standard periodic boundary [18]. One of the twisted boundaries preserves the $Z(3)$ invariance, which allows us to consider the spin at one end of the chain to differ by a factor $\omega^{\pm 1}$. The corresponding quantum spin chain is,
$H^{(\pm)}=-\frac{2}{\sqrt{3}}\sum_{j=1}^{L-1}\left(Z_{j}Z^{\dagger}_{j+1}+Z^{\dagger}_{j}Z_{j+1}+X_{j}+X^{\dagger}_{j}\right)-\frac{2}{\sqrt{3}}\left(\frac{Z_{L}Z^{\dagger}_{1}}{\omega^{\pm 1}}+\omega^{\pm 1}Z^{\dagger}_{L}Z_{1}+X_{L}+X^{\dagger}_{L}\right).$ (10)
It is natural to expect that the spectra of the above Hamiltonians should also be parametrized in terms of some transcendental set of Bethe ansatz type relations. As far as we know, this problem has not yet been considered in the literature, and in this paper, we will uncover the respective Bethe equation. We start by noticing that the spectra of $H^{(+)}$ and $H^{(-)}$ are related by complex conjugation, and therefore, we can concentrate our presentation to only one of the Hamiltonians. Here, we argue that the Bethe equation associated with the spectrum of the Hamiltonian $H^{(+)}$ is,
$\bigg[\frac{\sinh(\lambda_{j}+i\frac{\pi}{12})}{\sinh(\lambda_{j}-i\frac{\pi}{12})}\bigg]^{2L}=(-1)^{L}\exp\Big[{\frac{2i\pi Q}{3}}\Big]\prod_{\stackrel{{\scriptstyle k=1}}{{k\neq j}}}^{\bar{N}_{Q}}\frac{\sinh(\lambda_{j}-\lambda_{k}+i\frac{\pi}{3})}{\sinh(\lambda_{j}-\lambda_{k}-i\frac{\pi}{3})},~~j=1,\dots,\bar{N}_{Q},$ (11)
where we note the presence of an extra phase factor on the right-hand side of the Bethe equation when compared to the periodic case, see Eq.(6). In addition to that, numerical study of the spectrum for small lattice sizes indicates that the number of Bethe roots $\bar{N}_{Q}$ is now $\bar{N}_{0}=2L-2$ and $\bar{N}_{1}=\bar{N}_{2}=2L-1$.
If we denote by $E^{(+)}_{Q}$ the eigenvalues of the Hamiltonian $H^{(+)}$ we find that they can be expressed in terms of the solutions of the Bethe equation (11) as follows,
$E^{(+)}_{Q}=\sum_{j=1}^{\bar{N}_{Q}}\cot\Big(\frac{\pi}{12}-i\lambda_{j}\Big)+i\mu_{Q}-\frac{2L}{\sqrt{3}},$ (12)
where the sector dependent “chemical potential” factor $\mu_{Q}$ is given by
$\mu_{0}=0,~~\mu_{1}=-1,~~\mu_{2}=1.$ (13)
The second twisted boundary corresponds to taking the complex conjugation of the spin at the end of the toroidal chain. This boundary breaks the $Z(3)$ spin rotation and the quantum spin chain only commutes with the charge ${\cal{O}}_{Z(2)}$. The expression of the respective Hamiltonian is,
$H^{(c)}=-\frac{2}{\sqrt{3}}\sum_{j=1}^{L-1}\left(Z_{j}Z^{\dagger}_{j+1}+Z^{\dagger}_{j}Z_{j+1}+X_{j}+X^{\dagger}_{j}\right)-\frac{2}{\sqrt{3}}\left(Z_{L}Z_{1}+Z^{\dagger}_{L}Z^{\dagger}_{1}+X_{L}+X^{\dagger}_{L}\right)$ (14)
The Hilbert space of the Hamiltonian $H^{(c)}$ splits into two sectors associated with the distinct eigenvalues $\pm 1$ of the charge ${\cal{O}}_{Z(2)}$. For both sectors, we find that the corresponding Bethe equation is,
$\bigg[\frac{\sinh(\lambda_{j}+i\frac{\pi}{12})}{\sinh(\lambda_{j}-i\frac{\pi}{12})}\bigg]^{2L}=-(-1)^{L}\prod_{\stackrel{{\scriptstyle k=1}}{{k\neq j}}}^{2L}\frac{\sinh(\lambda_{j}-\lambda_{k}+i\frac{\pi}{3})}{\sinh(\lambda_{j}-\lambda_{k}-i\frac{\pi}{3})},~~j=1,\dots,2L,$ (15)
where we note the presence of an overall negative factor in Eq.(15) compared to the Bethe equation for the periodic case. In this case, the number of Bethe roots is fixed to $2L$ and the eigenvalues of the Hamiltonian $H^{(c)}$ are determined by,
$E^{(c)}=\sum_{j=1}^{2L}\cot\Big(\frac{\pi}{12}-i\lambda_{j}\Big)-\frac{2L}{\sqrt{3}}.$ (16)
We have organized this paper as follows. In the next section, we introduce toroidal boundary conditions for an integrable $n$-state spin model by exploring a recent mapping to an equivalent solvable $n$-state vertex model [19]. This permits us to formulate twisted boundary conditions in terms of boundary seams, responsible for changing the interactions at one end of the respective quantum spin chains. We apply this construction to the three-state scalar Potts model, making it possible to build the transfer matrices whose Hamiltonian limits give rise to the quantum spin chains $H^{(\pm)}$ and $H^{(c)}$ discussed above. In sections 3 and 4, we used certain properties of the transfer matrices eigenvalues to establish the Bethe equations parameterizing the spectra of the Hamiltonians $H^{(+)}$ and $H^{(c)}$. We investigate the solutions of the Bethe equations for $L=2$ and we use the Bethe roots to compute the Hamiltonian energies and the momenta of the eigenstates. In section 5, we use our framework to build integrable Hamiltonians with periodic boundary conditions whose spectra are obtained by combining the eigenvalues of the spin chains with different types of toroidal boundary conditions. In section 6, we present our concluding remarks. In the appendices $A$ and $B$, we summarize the analysis concerning the completeness of the solutions of the Bethe equations for $L=3$.
2 Toroidal Boundary Conditions
We start this section by describing the main results concerning an equivalence among integrable $n$-state spin and vertex models on the square lattice [19]. We recall that such a mapping assumes that the horizontal $W_{h}(i,j|x)$ and vertical $W_{v}(i,j|x)$ edge weights of the spin model are parametrized in terms of the difference of spectral parameters. It is well known that an important ingredient for the formulation of the integrability of a given vertex model is the corresponding Lax operator [20]. In this sense, it has been argued that such an operator can be constructed from the spin model edge weights as follows [19],
$\mathbb{L}_{12}(x)=\sum_{i,j,k=1}^{n}W_{h}(j,i|x)W_{v}(j,k|x)e_{ik}\otimes e_{ji},$ (17)
where $e_{i,j}$ denotes the standard $n\times n$ Weyl matrices.
In addition to that, it has been shown that one can exhibit a $\mathrm{R}$-matrix expressed also in terms of the spin edge weights,
$\mathrm{R}_{12}(x,y)=\sum_{i,j,k=1}^{n}\frac{W_{h}(j,i|x)W_{v}(j,k|x-y)}{W_{h}(k,i|y)}e_{ik}\otimes e_{ji},$ (18)
which together with the Lax operator satisfies the Yang-Baxter algebra [1, 20],
$\mathrm{R}_{12}(x,y)\mathbb{L}_{13}(x)\mathbb{L}_{23}(y)=\mathbb{L}_{23}(y)\mathbb{L}_{13}(x)\mathrm{R}_{12}(x,y).$ (19)
In this context, toroidal boundary conditions can be introduced by exploring certain internal symmetries of the Yang-Baxter algebra, see for instance [21]. They are associated with global transformations that leave the Yang-Baxter algebra invariant to preserve integrability. This is encoded by a group of matrices $\mathrm{G}$ satisfying the property,
$\mathrm{R}_{12}(x,y),\mathrm{G}\otimes\mathrm{G}]=0,$ (20)
An immediate consequence of this symmetry is the possibility to define a family of commuting transfer matrices as follows,
$\mathrm{T}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathbb{L}_{{\cal{A}}1}(x)\right],$ (21)
where the symbol $\cal{A}$ denotes the auxiliary space $\mathbb{C}^{n}$. We note that the periodic boundary condition corresponds to choosing $\mathrm{G}$ as the $n\times n$ identity matrix.
To define the corresponding quantum spin chain, we assume that the edge weights fulfill the following initial conditions,
$W_{h}(i,j|0)=1,~~~W_{v}(i,j|0)=\delta_{i,j},$ (22)
implying that at $x=0$ the Lax operator becomes the permutator $P_{12}$ on the tensor product $\mathbb{C}^{n}\otimes\mathbb{C}^{n}$, i.e. $\mathbb{L}_{12}(0)=\displaystyle\sum_{i,j=1}^{n}e_{ij}\otimes e_{ji}$. This condition also implies that the Lax operator satisfies property (20) since we have $\mathbb{L}_{12}(x)=\mathrm{R}_{12}(x,0)$.
As usual, the corresponding Hamiltonian is obtained by taking the logarithmic derivative of the transfer matrix at the point $x=0$ [1, 20]. Assuming that the boundary seam $\mathrm{G}$ is non-singular, we can write the Hamiltonian as,
$H=\sum_{j=1}^{L-1}\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}+\mathrm{H}^{(\mathrm{bound})}_{L,1}$, (23)
where the two-body terms $\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}$ and $\mathrm{H}^{(\mathrm{bound})}_{L1}$ are given by,
$\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}=P_{j,j+1}\frac{\mathrm{d}}{\mathrm{d}x}\mathbb{L}_{j,j+1}(x)\Big|_{x=0},~~~\mathrm{H}^{(\mathrm{bound})}_{L1}=\mathrm{G}_{L}^{-1}\mathrm{H}^{(\mathrm{bulk})}_{L,1}\mathrm{G}_{L}.$ (24)
At this point, we also note that $\mathrm{T}(0)$ plays the role of a translation operator which shifts the spin generators by one site modulo $L$. More precisely, one can verify the following relations,
$\mathrm{T}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}\right]\mathrm{T}^{-1}(0)=\mathrm{H}^{(\mathrm{bulk})}_{j+1,j+2},~~j=1,\dots,L-2,$
$\mathrm{T}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{L-1,L}\right]\mathrm{T}^{-1}(0)=\mathrm{H}^{(\mathrm{bound})}_{L,1}$. (25)
Let us now discuss the above construction in the specific case of the three-state scalar Potts model. The first step is to solve Eq.(20) for the $\mathrm{R}$-matrix (18) of the equivalent three-state vertex model. Starting with a generic $3\times 3$ $\mathrm{G}$ matrix and considering the structure of the edge weights (1) we find that there exist three basic solutions besides the identity matrix. It turns out that any non-trivial solution to Eq.(20) can be seen as a composition of the following matrices,
$\mathrm{G}^{(+)}={X}^{\dagger},~~\mathrm{G}^{(-)}=X,~~\mathrm{G}^{(c)}=C$ (26)
and as a result, we can define three families of commuting transfer matrices,
$\mathrm{T}^{(\pm)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(\pm)}_{{\cal A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right],~~~\mathrm{T}^{(c)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(c)}_{\cal{A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right].$ (27)
The transfer matrices $\mathrm{T}^{(\pm)}(x)$ are invariant under the $Z(3)$ spin rotation and their Hamiltonian limits, apart from a constant bulk term, correspond to the quantum spin chains $H^{(\pm)}$. Thus, we have the following commutation relations,
$\mathrm{T}^{(\pm)}(x),{\cal{O}}_{Z(3)}]=0,~~[\mathrm{T}^{(\pm)}(x),H^{(\pm)}]=0$ (28)
such that the spectra of $\mathrm{T}^{(+)}(x)$ and $\mathrm{T}^{(-)}(x)$ are related by the following equivalences,
$\mathrm{Spec}\big[T^{(+)}\big]_{Q=0}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=0},~~\mathrm{Spec}\big[T^{(+)}\big]_{Q=1}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=2},~~\mathrm{Spec}\big[T^{(+)}\big]_{Q=2}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=1}$ (29)
The transfer matrix $\mathrm{T}^{(c)}(x)$ keeps only the invariant concerning the $Z(2)$ subgroup associated with the charge conjugation. The respective Hamiltonian limit gives rise to the quantum spin chain $H^{(c)}$, namely
$\mathrm{T}^{(c)}(x),{\cal{O}}_{Z(2)}]=0,~~[\mathrm{T}^{(c)}(x),H^{(c)}]=0$ (30)
In the next sections, we shall see that the construction of these transfer matrices is essential to set up the Bethe equations for the corresponding quantum spin chains. From the spectral relations (29) it is enough to consider the problem for $T^{(+)}(x)$ and $T^{(c)}(x)$.
3 The Bethe equation for $T^{(+)}(x)$
Here, we follow the strategy devised by Baxter to solve the eigenvalue problem associated with the diagonal-to-diagonal transfer matrix of the Ising model [1]. We recall that this kind of approach has also been used to uncover the Bethe equation for the critical $Z(n)$ spin model with periodic boundary conditions [11, 22]. The first step is to build a suitable proposal for the form of the eigenvalues of the transfer matrix $T^{(+)}$. To this end, we observe any matrix elements of such operator have the following structure,
$\left[T^{(+)}\right]_{a_{1},\dots,a_{L}}^{b_{1},\dots,b_{L}}=\Big[\frac{1}{g(x)g_{1}(x)}\Big]^{L}H(x)$ (31)
where $g(x)$ and $g_{1}(x)$ take into account the denominators of the edge weights,
$g(x)=\sin(\frac{\pi}{6}+x),~~g_{1}(x)=\sin(\frac{\pi}{3}-x)$ (32)
After the denominators have been removed, each entry of the transfer matrix is expressed in terms of the product of sines. Therefore, the polynomial $H(x)$ can be factorized with the help of exponentials and as a result it can be written as,
$H(x)=\prod_{j=1}^{2L}\Big(c_{j}\exp(ix)+d_{j}\exp(-ix)\Big)$ (33)
where $c_{j}$ and $d_{j}$ are coefficients independent of $x$.
Considering that the transfer matrix commutes for distinct values of the spectral parameter, we can write the eigenvalue problem as follows,
$T^{(+)}(x)\left|\phi\right\rangle=\Lambda^{(+)}(x)\left|\phi\right\rangle$ (34)
where the eigenstates $\left|\phi\right\rangle$ do not depend on $x$.
Due to the independence of the eigenstates on $x$ a given eigenvalue $\Lambda^{(+)}(x)$ can be regarded as a linear combination of the elements of the transfer matrix $T^{(+)}(x)$ with constant coefficients. As a consequence, we can conclude that $\Lambda^{(+)}(x)$ should have the same form as given by Eqs.(31,33). To obtain further insights on the structure of $\Lambda^{(+)}(x)$ we have studied the eigenvalue problem (34) for small finite sizes. This analysis reveals to us that the the form of eigenvalues depends on the $Z(3)$ sectors besides not given solely in terms of the product of sines. Investigating the asymptotic behavior of $\Lambda^{(+)}(x)$ for $x\rightarrow\pm i\infty$ leads to propose the following form,
$\Lambda^{(+)}(x)=\varrho\Big[\frac{1}{g(x)g_{1}(x)}\Big]^{L}\exp\Big[i\mu_{Q}(\frac{\pi}{6}-x)\Big]\prod_{k=1}^{\bar{N}_{Q}}\sin(\xi_{k}-\frac{\pi}{6}+x)$ (35)
where $\varrho$ is a normalization, the factor $\mu_{Q}$ is given by Eq.(13) and the variables $\xi_{k}$ are related to the zeros of the eigenvalues, yet to be determined. At this stage of the analysis, the number $\bar{N}_{Q}$ is bounded by $2L$. We remark that the above structure for the eigenvalue is similar to that used for periodic boundary conditions [11] except by the presence of the exponential factor in Eq.(35).
It turns out that the normalization can be fixed by observing that the transfer matrix at $x=\frac{\pi}{6}$ reduces to the identity matrix. As a consequence of that, the final form for the eigenvalue is,
$\Lambda^{(+)}(x)=\Big[\frac{g(\frac{\pi}{6})g_{1}(\frac{\pi}{6})}{g(x)g_{1}(x)}\Big]^{L}\exp\Big[i\mu_{Q}(\frac{\pi}{6}-x)\Big]\prod_{k=1}^{\bar{N}_{Q}}\frac{\sin(\xi_{k}-\frac{\pi}{6}+x)}{\sin(\xi_{k})}$ (36)
and in this normalized form, we note that the exponential factor can be formally interpreted as the presence of a zero $\xi_{k}$ at $\pm i\infty$ for arbitrary $L$. However, the explicit computation of the eigenvalues requires working with variables defined on a finite interval, and in practice, we have to split potential asymptotic contributions from the standard ratio of sines.
The next step is to determine the set of nonlinear equations satisfied by the variables $\xi_{k}$ on a finite interval. This task can be done by exploring certain matrix identity among the products of transfer matrices with suitable distinct spectral parameters. Here we have been inspired by the matrix structure derived for the diagonal-to-diagonal transfer matrix of the three-state Chiral Potts model [12, 13]. We have found that the transfer matrix $T^{(+)}(x)$ satisfies the following identity,
$T^{(+)}(x-\frac{\pi}{3})T^{(+)}(x-\frac{\pi}{6})T^{(+)}(x)$ $=$ $T^{(+)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T^{(+)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T^{(+)}(x)$ (37)
$+$ $\big[f_{3}(x)\big]^{L}T^{(+)}(x+\frac{\pi}{3})\bigg\}$
where the auxiliary functions $f_{1}(x)$, $f_{2}(x)$ and $f_{3}(x)$ are determined by,
$f_{1}(x)=3\tan(x)\cot(x+\frac{\pi}{6}),~~f_{2}(x)=3\tan(x-\frac{\pi}{6})\cot(x),~~f_{3}(x)=3\tan(x-\frac{\pi}{6})\cot(x+\frac{\pi}{6})$ (38)
We can turn the above matrix equation into a functional relation for the eigenvalues by acting on both sides of Eq.(37) on a common set of eigenvectors. In this procedure we have used the fact that the transfer matrices with different spectral parameters commute. If we choose $x$ to be any of the zeros of a given eigenvalue, we see that the left-hand side of Eq.(37) vanishes when it acts on the respective eigenvector. Now by taking $x=\frac{\pi}{6}-\xi_{j}$, the right hand side of Eq.(37) produces the following condition for the eigenvalues,
$f_{1}(\frac{\pi}{6}-\xi_{j})\big]^{L}\Lambda^{(+)}(-\frac{\pi}{6}-\xi_{j})+[f_{3}(\frac{\pi}{6}-\xi_{j})\big]^{L}\Lambda^{(+)}(\frac{\pi}{2}-\xi_{j})=0$ (39)
To obtain the constraints for the variables $\xi_{k}$ we substitute in Eq.(39) the expressions for the eigenvalue (36) and for the auxiliary functions (38). After some simplifications we find that the Bethe equation for these variables are given by,
$\bigg[\frac{\sin(\xi_{j}-\frac{\pi}{6})}{\sin(\xi_{j})}\bigg]^{2L}=(-1)^{L}\exp\Big[{\frac{-2i\pi\mu_{Q}}{3}}\Big]\prod_{\stackrel{{\scriptstyle k=1}}{{k\neq j}}}^{\bar{N}_{Q}}\frac{\sin(\xi_{k}-\xi_{j}+\frac{\pi}{3})}{\sin(\xi_{k}-\xi_{j}-\frac{\pi}{3})},~~j=1,\dots,\bar{N}_{Q}$ (40)
We next recall that the energy of the Hamiltonian $H^{(+)}$ is computed by taking the first derivative of the logarithm of the eigenvalue at the value $x=0$. Considering the normalization we have used for the edge weights we find,
$E_{Q}^{(+)}=-\frac{1}{\Lambda^{(+)}(0)}\frac{\mathrm{d}}{\mathrm{d}x}\Lambda^{(+)}(x)\Big|_{x=0}-4L\cot(\frac{\pi}{3})=\sum_{j=1}^{\bar{N}_{Q}}\cot\Big(\frac{\pi}{6}-\xi_{j}\Big)+i\mu_{Q}-\frac{2L}{\sqrt{3}}$ (41)
At this point, we note that Eqs.(40,41) leads us to the expressions of the Bethe equation (11) and the energy eigenvalue (12) presented in the introduction by choosing,
$\xi_{j}=i\lambda_{j}+\frac{\pi}{12}$ (42)
[…]
$s_{p}=\frac{iL}{2\pi}\ln\big[\Lambda^{(+)}(0)\big]$ $=$ $\frac{iL}{2\pi}\sum_{k=1}^{\bar{N}_{Q}}\ln\bigg[\frac{\sin(\xi_{k}-\frac{\pi}{6})}{\sin(\xi_{k})}\bigg]-\frac{L}{12}\mu_{Q}$ (43)
$=$ $\frac{iL}{2\pi}\sum_{k=1}^{\bar{N}_{Q}}\ln\bigg[\frac{\sinh(\lambda_{k}+i\frac{\pi}{12})}{\sinh(\lambda_{k}-i\frac{\pi}{12})}\bigg]-\frac{L}{12}\mu_{Q}$
3.1 The Bethe roots for $L=2$
We now describe the set of roots solving the Bethe equation (11) associated with the energy eigenvalues of the Hamiltonian $H^{(+)}$ for $L=2$. We obtain some insight on the nature of the roots, exploring the fact that the eigenvectors of $T^{(+)}(x)$ are independent of the spectral parameter. By choosing a suitable value for $x$ we can compute numerically the eigenvectors making it possible the determination of the corresponding eigenvalues as polynomial functions on $x$. The zeros of such polynomials work as starting values for the numerical solution of the Bethe equation (11) providing us the Bethe roots with high precision.
In table (1) we have summarized our results for $L=2$ considering the total number of nine eigenstates. For each energy of the spectrum of $H^{(+)}$ we exhibit the sector $Q$, the spin $s_{p}$ of the state and the corresponding roots $\{\lambda_{j}\}$ solving the Bethe equation (11). The energy values are obtained with the help of Eqs.(12,13) and we have verified that they agree with those obtained from the exact diagonalization of the Hamiltonian $H^{(+)}$. From table (1) we note that if a given energy in sector $Q=1$ is parametrized by a set of roots $\{\lambda_{j}\}$ the same energy is obtained in sector $Q=2$ using the opposite set of roots $\{-\lambda_{j}\}$. From our solution, we see that this property is valid for arbitrary $L$ in accordance with the fact that the spectra for sectors $Q=1,2$ are the same. In addition, from Eqs.(13,43) we see that the momenta of such states are however opposite to each other.
$E^{(+)}$ $Q$ $s_{p}$ $\lambda_{j}$
-4.93624921 0 0 $\pm 0.53202156i$
-2.30940107 1 $\frac{1}{3}$ $0.06315064\pm 0.54254935i,~0.19822180$
-2.30940107 2 $\frac{1}{3}$ $0.06315064\pm 0.54254935i$0.19822180-0.19822180
-0.84529946 1 $\frac{2}{3}$ $0.16532518\pm 0.50188873i$0.10277717-0.10277717
-0.84529946 2 $\frac{2}{3}$ $0.16532518\pm 0.50188873i,~0.10277717$
1.15470053 0 1 $0,~\frac{i\pi}{2}$
3.15470053 1 $\frac{2}{3}$ 0.245380170.245380170.02638725-0.026387250.64959867-0.64959867
3.15470053 2 $\frac{2}{3}$ 0.24538017-0.245380170.02638725,0.649598670.02638725,~0.64959867
3.78154867 0 0 $\pm 0.12258177$
Table 1: The eigenvalues ($E^{(+)}$) of the Hamiltonian $H^{(+)}$ (10) for $L=2$. For each energy, we indicate the sector ($Q$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$).
Inspecting table (1) we see the occurrence of low-lying eigenstates whose values for the spin are $\pm\frac{1}{3}$ and $\pm\frac{2}{3}$. We shall argue that this feature is in accordance with the operator content proposed for the critical three-state Potts model with a $Z(3)$ invariant twisted boundary condition [17, 18]. We first recall that the primary fields present in this critical system determine the asymptotic finite-size behaviour of the energies and the momenta of the low-lying excitations [23, 24]. For a given primary field with scaling dimension $X=\Delta+\bar{\Delta}$ and spin $S=\Delta-\bar{\Delta}$ we have,
$E(L)=e_{\infty}L+\frac{2\pi v_{F}}{L}\big(\Delta+\bar{\Delta}-\frac{c}{12}+n_{1}+n_{2}\big)+{\mathcal{O}}(1/L),$
$P(L)=\frac{2\pi}{L}\big(\Delta-\bar{\Delta}+n_{1}-n_{2}\big),~~n_{1},n_{2}=0,1,2,\dots,$ (44)
where $e_{\infty}$ is the ground state energy in the thermodynamic limit and $v_{F}$ is the Fermi velocity of the low-lying excitations [11, 15]. It turns out that for the $Z(3)$ twisted three-state Potts model, the corresponding conformal weights $(\Delta,\bar{\Delta})$ are [17, 18],
$(\Delta,\bar{\Delta})\big|_{Q=0}=\big(\frac{1}{15},\frac{1}{15}\big),~\big(\frac{2}{3},\frac{2}{3}\big)$
$(\Delta,\bar{\Delta})\big|_{Q=1}=\big(\frac{1}{15},\frac{2}{5}\big),~\big(\frac{2}{3},0\big),~\big(\frac{1}{15},\frac{7}{5}\big),~\big(\frac{2}{3},3\big)$
$(\Delta,\bar{\Delta})\big|_{Q=2}=\big(\frac{2}{5},\frac{1}{15}\big),~\big(0,\frac{2}{3}\big),~\big(\frac{7}{5},\frac{1}{15}\big),~\big(3,\frac{2}{3}\big)$ (45)
From this operator content, we see that second lowest energy state is expected to be associated with $X=\frac{7}{15}$ while the third energy state should have $X=\frac{2}{3}$, being both doubly degenerate states since they occur in the sectors $Q=1,2$. The conformal weights associated with the second scaling dimension are $(\frac{1}{15},\frac{2}{5})\oplus(\frac{2}{5},\frac{1}{15})$ and the conformal spins of such states are therefore $S=\mp\frac{1}{3}$. For the third scaling dimension, the conformal weights are $(\frac{2}{3},0)\oplus(0,\frac{2}{3})$ and hence these states have conformal spins $S=\pm\frac{2}{3}$. Considering table (1) we conclude that such conformal spins indeed coincide with the values $s_{p}$ obtained from the Bethe roots for the second and the third double degenerate energies of the Hamiltonian spectrum.
We have performed the above finite size computations for $L=3$ and the results are presented in Appendix B. This analysis confirms the conclusions we have made for the momenta of the states discussed above. We have also investigated the nature of the Bethe roots for several low-lying states with lattice size $4\leq L\leq 7$ and our results indicate that the number of the Bethe roots for each sector should be,
$\bar{N}_{0}=2L-2,~~\bar{N}_{1}=\bar{N}_{2}=2L-1$ (46)
4 The Bethe equation for $T^{(c)}(x)$
Here, we use the same procedure explained in the previous section to determine the parameterization of the spectrum of Hamiltonian $H^{(c)}$ in terms of a Bethe ansatz equation. We therefore avoid the repetition of technical details and concentrate our presentation to the main results. In particular, we find out that the structure of the eigenvalues of $T^{(c)}(x)$ can be given as follows,
$\Lambda^{(c)}(x)=\Big[\frac{g(\frac{\pi}{6})g_{1}(\frac{\pi}{6})}{g(x)g_{1}(x)}\Big]^{L}\prod_{k=1}^{2L}\frac{\sin(\xi_{k}-\frac{\pi}{6}+x)}{\sin(\xi_{k})}$ (47)
where now the number of sine terms is the same as that present in the general form (33) for the transfer matrix elements. We arrive at this conclusion by inspecting the asymptotic limit of the eigenvalues of $T^{(c)}(x)$ at the points $\pm i\infty$ for small lattice sizes.
As before, the Bethe equation for the variables $\xi_{k}$ is uncovered with the help of a functional relation for the eigenvalue involving distinct spectral parameters. This equation is derived from a matrix identity for the product of transfer matrices whose main structure is similar to that presented in previous section, see Eq.(37). More specifically, we find that $T^{(c)}(x)$ satisfies the following matrix identity,
$T^{(c)}(x-\frac{\pi}{3})T^{(c)}(x-\frac{\pi}{6})T^{(c)}(x)$ $=$ $T^{(c)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T^{(c)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T^{(c)}(x)$ (48)
- $\big[f_{3}(x)\big]^{L}T^{(c)}(x+\frac{\pi}{3})\bigg\}$
where the main difference with Eq.(37) is the presence of a minus sign on the term proportional to the amplitude $f_{3}(x)$.
If we choose $x=\frac{\pi}{6}-\xi_{k}$ the above matrix identity can turn into a constraint for the eigenvalues (47) producing a set of relations for the variables $\xi_{k}$. We find that such a Bethe equation is given by,
$\bigg[\frac{\sin(\xi_{j}-\frac{\pi}{6})}{\sin(\xi_{j})}\bigg]^{2L}=(-1)^{L+1}\prod_{\stackrel{{\scriptstyle k=1}}{{k\neq j}}}^{2L}\frac{\sin(\xi_{k}-\xi_{j}+\frac{\pi}{3})}{\sin(\xi_{k}-\xi_{j}-\frac{\pi}{3})},~~j=1,\dots,2L$ (49)
which becomes Eq.(15) when we once again set $\xi_{j}=i\lambda_{j}+\frac{\pi}{12}$.
From the expression for the eigenvalue (47), we can compute both the energy and the momenta of a given state. The energy $E^{(c)}$ is given by,
$E^{(c)}=-\frac{1}{\Lambda^{(c)}(0)}\frac{\mathrm{d}}{\mathrm{d}x}\Lambda^{(c)}(x)\Big|_{x=0}-4L\cot(\frac{\pi}{3})$ $=$ $\sum_{j=1}^{2L}\cot\Big(\frac{\pi}{6}-\xi_{j}\Big)-\frac{2L}{\sqrt{3}}$ (50)
$=$ $\sum_{j=1}^{2L}\cot\Big(\frac{\pi}{12}-i\lambda_{j}\Big)-\frac{2L}{\sqrt{3}},$
while the spin $s_{p}$ of the eigenstate is,
$s_{p}=\frac{iL}{2\pi}\ln\big[\Lambda^{(c)}(0)\big]=\frac{iL}{2\pi}\sum_{k=1}^{2L}\ln\bigg[\frac{\sin(\xi_{k}-\frac{\pi}{6})}{\sin(\xi_{k})}\bigg]=\frac{iL}{2\pi}\sum_{k=1}^{2L}\ln\bigg[\frac{\sinh(\lambda_{k}+i\frac{\pi}{12})}{\sinh(\lambda_{k}-i\frac{\pi}{12})}\bigg].$ (51)
4.1 Bethe roots for L=2
We now proceed to investigate the Bethe equation parameterization we have found for the spectrum of the Hamiltonian $H^{(c)}$ with $L=2$. This Hamiltonian has a $Z(2)$ invariance and the Hilbert space splits into two sectors labeled by the eigenvalues of the charge ${\cal{O}}_{Z(2)}$. If we denote these eigenvalues by the symbol $\nu_{c}$ we find that there are $(3^{L}+1)/2$ states with $\nu_{c}=+1$ and $(3^{L}-1)/2$ states with $\nu_{c}=-1$. We obtain the initial values for the Bethe roots by computing the eigenvalues of the transfer matrix $T^{(c)}(x)$ as meromorphic functions on $x$. We use this data to solve the Bethe equation (15) with high numerical precision and the energy and the spin of the states are calculated with the help of Eqs.(16,51). The energies obtained from the Bethe roots are then checked with those coming from the exact diagonalization of the Hamiltonian $H^{(c)}$. The results are shown in table (2).
$E^{(c)}$ $\nu_{c}$ $p$ $\lambda_{j}$
-5.77350269 +1 0 $0.13588376\pm 0.57005210i$$0.13588376\pm 0.57005210i$
-3.89397505 +1 0 $\pm 0.50828016i,~\pm 0.524244017$
[…]
1.154700538 +1 11 $0.28844398\pm 0.61762004i$0.558766300.558766300.018121660.01812166
1.154700538 +1 1-1 $0.28844398\pm 0.61762004i$0.55876630-0.558766300.01812166-0.01812166
2.739274521 +1 0 $\pm 1.08877871i$$\pm 0.14111440$
[…]
Table 2: The eigenvalues ($E^{(c)}$) of the Hamiltonian $H^{(c)}$ (14) for $L=2$. For each energy, we indicate the sector ($\nu_{c}$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$).
Interestingly enough, we observe that there are excitations with spin $\pm\frac{1}{2}$ in the charge sector with negative eigenvalue. This is in accordance with the operator content predicted for the three-state Potts model with a $Z(2)$ twisted boundary condition. In this case, the expected conformal weights are [18],
$(\Delta,\bar{\Delta})\big|_{\nu_{c}=+1}=\big(\frac{1}{40},\frac{1}{40}\big),~\big(\frac{1}{8},\frac{1}{8}\big),~\big(\frac{21}{40},\frac{21}{40}\big),~\big(\frac{13}{8},\frac{13}{8}\big)$
$(\Delta,\bar{\Delta})\big|_{\nu_{c}=-1}=\big(\frac{21}{40},\frac{1}{40}\big),~\big(\frac{1}{40},\frac{21}{40}\big),~\big(\frac{13}{8},\frac{1}{8}\big),~\big(\frac{13}{8},\frac{1}{8}\big).$ (52)
Considering this operator content, we note that the third lowest energy state is expected to have scaling dimension $X=\frac{11}{20}$. The respective conformal weights are $(\frac{21}{40},\frac{1}{40})\oplus(\frac{1}{40},\frac{21}{40})$ and this means that the conformal spins are $\pm\frac{1}{2}$. From table (2) we see that these are exactly the spin values of the double degenerate third energy eigenstate of the Hamiltonian occurring in the sector $\nu_{c}=-1$. In appendix $C$ we have presented the Bethe roots data concerning the completeness of the spectrum for $L=3$. Once again, we observe that such low-lying excitation has half-integer spin, and this feature has also been checked up to $L=7$.
5 Other families of integrable Hamiltonians
We can explore the invariance of the Yang-Baxter algebra (20) to construct more general families of commuting transfer matrices. For example, we can write,
$\mathrm{\tilde{T}}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}1}(x)\right],$ (53)
where the boundary seams are spread throughout the bulk by acting them on each of the Lax operators.
The Hamiltonian of the corresponding quantum spin chain is obtained as a logarithmic derivative of the transfer matrix at point $x=0$. As a result, we have,
$\tilde{H}=\sum_{j=1}^{L}\mathrm{G}^{-1}_{j}\Bigg[P_{j,j+1}\frac{\mathrm{d}}{\mathrm{d}x}\mathbb{L}_{j,j+1}(x)\Big|_{x=0}\Bigg]\mathrm{G}_{j}$ (54)
where the periodic boundary condition $L+1\equiv 1$ is assumed.
Let us apply the above construction for the classical three-state scalar Potts spin model. If we take the boundary matrix $\mathrm{G}=X^{\dagger}$ one finds that the Hamiltonian is given by,
$\tilde{H}_{1}=-\frac{2}{\sqrt{3}}\sum_{j=1}^{L}\left(\frac{1}{\omega}Z_{j}Z^{\dagger}_{j+1}+\omega Z^{\dagger}_{j}Z_{j+1}+X_{j}+X^{\dagger}_{j}\right).$ (55)
It turns out that the bulk couplings $\omega^{\pm 1}$ in the Hamiltonian (55) can be transformed into the twisted boundary conditions depending on the size of the lattice. In fact, by performing the following local similarity transformation,
$Z_{j}\rightarrow\omega^{-j}Z_{j},~~Z^{\dagger}_{j}\rightarrow\omega^{j}Z^{\dagger}_{j},~~X_{j}\rightarrow X_{j},$ (56)
the Hamiltonian (55) takes the form,
$\tilde{H}_{1}=-\frac{2}{\sqrt{3}}\sum_{j=1}^{L-1}\left(Z_{j}Z^{\dagger}_{j+1}+Z^{\dagger}_{j}Z_{j+1}+X_{j}+X^{\dagger}_{j}\right)-\frac{2}{\sqrt{3}}\left(\frac{Z_{L}Z^{\dagger}_{1}}{\omega^{L}}+\omega^{L}Z^{\dagger}_{L}Z_{1}+X_{L}+X^{\dagger}_{L}\right).$ (57)
We note that the boundary term of the Hamiltonian (57) depends on the lattice size modulo 3. For $L=3m$, where $m$ is a positive integer, the twisted term becomes unity, and we have periodic boundary conditions, but when $L=3m\pm 1$ we have twisted boundary conditions associated with the matrices $\mathrm{G}^{(\pm)}$. Therefore, the spectrum of the Hamiltonian (55) can be obtained from the spectral relations,
$\mathrm{Spec}\Big[\tilde{H}_{1}\Big]_{L=3m}=\mathrm{Spec}\Big[H^{(p)}\Big]_{L=3m},~~\mathrm{Spec}\Big[\tilde{H}_{1}\Big]_{L=3m\pm 1}=\mathrm{Spec}\Big[H^{(\pm)}\Big]_{L=3m\pm 1}.$ (58)
We now consider the situation in which the boundary matrix is $\mathrm{G}=C$. In this case the corresponding Hamiltonian is,
$\tilde{H}_{2}=-\frac{2}{\sqrt{3}}\sum_{j=1}^{L}\left(Z_{j}Z_{j+1}+Z^{\dagger}_{j}Z^{\dagger}_{j+1}+X_{j}+X^{\dagger}_{j}\right).$ (59)
Once again, the eigenvalues of the Hamiltonian (59) can be determined in terms of the spectrum of the three-state Potts quantum chain with boundary conditions depending on the lattice size. This can be seen by considering the charge conjugation symmetry only for even sites, namely
$Z_{j}\rightarrow Z^{\dagger}_{j},~~Z^{\dagger}_{j}\rightarrow Z_{j},~~X_{j}\rightarrow X^{\dagger}_{j},~~X^{\dagger}_{j}\rightarrow X_{j}~~\mathrm{for}~~j=2,4,6,\dots.$ (60)
By applying this transformation, we find that the Hamiltonian (59) changes to,
$\tilde{H}_{2}=-\frac{2}{\sqrt{3}}\sum_{j=1}^{L-1}\left(Z_{j}Z^{\dagger}_{j+1}+Z^{\dagger}_{j}Z_{j+1}+X_{j}+X^{\dagger}_{j}\right)+\tilde{H}^{(\mathrm{bound})}_{2}$ (61)
where the boundary term $\tilde{H}^{(bound)}_{2}$ depends on the parity of the lattice size,
$\tilde{H}^{(\mathrm{bound})}_{2}=\Bigg\{\begin{array}[]{c}-\frac{2}{\sqrt{3}}\left(Z_{L}Z^{\dagger}_{1}+Z^{\dagger}_{L}Z_{1}+X_{L}+X^{\dagger}_{L}\right)~~\mathrm{for}~~L\in\mathrm{even}\\ -\frac{2}{\sqrt{3}}\left(Z_{L}Z_{1}+Z^{\dagger}_{L}Z^{\dagger}_{1}+X_{L}+X^{\dagger}_{L}\right)~~\mathrm{for}~~L\in\mathrm{odd},\end{array}$ (62)
and as a result, we have the following spectral equivalence,
$\mathrm{Spec}\Big[\tilde{H}_{2}\Big]_{L=\mathrm{even}}=\mathrm{Spec}\Big[H^{(p)}\Big]_{L=\mathrm{even}},~~\mathrm{Spec}\Big[\tilde{H}_{2}\Big]_{L=\mathrm{odd}}=\mathrm{Spec}\Big[H^{(c)}\Big]_{L=\mathrm{odd}}.$ (63)
The quantum spin chain (59) appears to be an interesting integrable variant of the three-state Potts spin chain with periodic boundary conditions. A nice feature of such spin chain is that it provides a lattice realization of the complete set of the lowest weights of the Virasoro minimal model with central charge $c=4/5$. To this end let us denote by $\phi_{r,s}$ the primary fields of this theory whose conformal weights are given by [16],
$\Delta_{r,s}=\frac{(6r-5s)^{2}-1}{120},~~1\leq r\leq 2,~~1\leq s\leq 5$ (64)
where the underlying operator algebra is invariant under a $Z(2)$ invariance of the form [25],
$\phi_{r,s}\rightarrow(-1)^{s+1}\phi_{r,s}$ (65)
For $L$ even the spectrum of $\tilde{H}_{2}$ is the same of the three-state Potts spin chain with periodic boundary conditions whose operator content was discussed in [16, 17]. The underlying conformal weights correspond to that of the primary fields which are even under the $Z(2)$ invariance (65),
$\Delta_{1,3}=\frac{2}{3},~\Delta_{1,5}=3,~\Delta_{2,1}=\frac{2}{5},~\Delta_{2,3}=\frac{1}{15},~\Delta_{2,5}=\frac{7}{5}$ (66)
but when $L$ is odd the spectrum of $\tilde{H}_{2}$ coincides with that of the three-state Potts spin chain with a $Z(2)$ twisted boundary. The operator content was proposed in [18] and the respective conformal weights correspond to those of the primary fields which are odd under the $Z(2)$ invariance (65),
$\Delta_{1,2}=\frac{1}{8},~\Delta_{1,4}=\frac{13}{8},~\Delta_{2,2}=\frac{1}{40},~\Delta_{2,4}=\frac{21}{40}$ (67)
and altogether give the values of the lowest weights predicted by Eq.(64).
6 Conclusions
The main goal of this paper was to establish the Bethe type equations for the spectra of the critical three-state Potts quantum spin chain with twisted boundary conditions. To accomplish this task, we first constructed the respective transfer matrices exploring a mapping among integrable classical spin and vertex models. This equivalence permits us to formulate a toroidal boundary condition in terms of a boundary seam responsible for changing the interaction among the spins at one end of the chain. Exploring the internal symmetries of the Yang-Baxter algebra we find that there are two other classes of integrable twisted boundaries besides the periodic boundary condition.
The first class of twisted boundary preserves the $Z(3)$ invariance of the three-state Potts model. In this case, the ansatz for the transfer matrix eigenvalue needs a suitable modification compared with that devised for the periodic boundary condition [11]. The second class of twisted boundary keeps only the $Z(2)$ invariance and the transfer matrix eigenvalue has the same form of the periodic case, but now the number of the Bethe roots is fixed. In both cases, the corresponding Bethe equations are found by exploring certain matrix identities for the product of transfer matrices. The structure of these relations are similar for both twisted boundaries except for a sign in one of the amplitudes of the linear combination of transfer matrices. The basic form of the Bethe equations is that found for the periodic boundary condition except by the presence of phase factors directly related to the ratio of the eigenvalues of the matrices representing the boundary seams. In this sense, the situation is equivalent to that found for integrable vertex systems invariant under continuous symmetries, such as the six-vertex model [1, 26]. We have studied all the eigenvalues of the twisted three-state Potts spin chain for $L=2,3$ in terms of the root content of the corresponding Bethe equations. This allowed us to compute the momenta of the eigenstates and we found that some of the low-lying excitations have fractional values for the respective spin. We have argued that they agree with the conformal spins predicted by the corresponding conformal field theories.
It is conceivable that the framework discussed in this paper may be applied to other integrable classical spin model, such as those invariant under the $Z(n)$ group of discrete rotations in the spin space. Recall here that the $Z(n)$ algebra is defined by,
$Z^{n}=X^{n}=1,~~ZX=\omega XZ,~~\omega=\exp(2\pi i/n)$ (68)
where the $Z$ and $X$ are $n\times n$ operators generalizing the Pauli matrices,
$Z=\left(\begin{array}[]{cccccc}1&0&0&\dots&0&0\\ 0&\omega&0&\dots&0&0\\ 0&0&\omega^{2}&&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\dots&\omega^{n-2}&0\\ 0&0&0&\dots&0&\omega^{n-1}\\ \end{array}\right)_{n\times n}~~~X=\left(\begin{array}[]{cccccc}0&0&0&\dots&0&1\\ 1&0&0&\dots&0&0\\ 0&1&0&&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\dots&0&0\\ 0&0&0&\dots&1&0\\ \end{array}\right)_{n\times n}$~~~ (69)
The typical example is the Fateev-Zamolodchikov model whose edge weights are given by [5],
$W_{h}(a,b|x)=\prod_{j=1}^{|a-b|}\frac{\sin\left((2j-1)\frac{\pi}{2n}-x\right)}{\sin\left((2j-1)\frac{\pi}{2n}+x\right)},~~W_{v}(a,b|x)=\prod_{j=1}^{|a-b|}\frac{\sin\left((j-1)\frac{\pi}{n}+x\right)}{\sin\left(\frac{j\pi}{n}-x\right)},$ (70)
and the associated quantum spin chain with periodic boundary conditions is given by [8, 9],
$H_{n}^{(p)}=-\sum_{j=1}^{L}\sum_{k=1}^{n}\frac{1}{\sin(k\pi/n)}\Big(\left[Z_{j}Z^{\dagger}_{j+1}\right]^{k}+\left[X_{j}\right]^{k}\Big),~~Z^{\dagger}_{L+1}=Z^{\dagger}_{1}$ (71)
We have investigated the symmetries of the Yang-Baxter algebra associated with the equivalent vertex model built from the edge weights of the Fateev-Zamolodchikov model. We find that there are (n1)(n-1) integrable toroidal boundary conditions compatible with the $Z(n)$ invariance besides the above mentioned periodic boundary condition. The respective boundary seams are given by,
$\mathrm{G}^{(l)}=X^{n-l},~~l=1,\dots,n-1$ (72)
which now leads us to the following twisted boundary conditions,
$Z^{\dagger}_{L+1}=\omega^{-l}Z^{\dagger}_{1}.$ (73)
For $n\geq 3$, we also have an integrable twisted boundary condition associated with the breaking of the $Z(n)$ rotational invariance of the spin model. In this case the boundary seam is,
[…]
and the respective toroidal boundary condition corresponds to charge conjugation at the end of the chain,
$Z^{\dagger}_{L+1}=Z_{1}.$ (75)
In principle, it should be possible to derive the Bethe equations for the $Z(n)$ Fateev-Zamolodichikov model with the above toroidal boundary conditions. For the boundary seams preserving the $Z(n)$ invariance, we expect that the expression for the eigenvalue of the transfer matrix should contain additional suitable exponential terms that multiply the standard product of sine ratios. At present, however, the precise determination of the dependence of the respective phase factors on the many possible $Z(n)$ sectors is an open question. Another relevant technical issue we have to deal with is the derivation of the functional relations constraining the eigenvalues for arbitrary spectral parameters. In any case, one expects that the structure of the Bethe equations will be dependent on whether $n$ is odd or even since this is the situation for periodic boundary condition [22]. We hope to return to some of these problems in future work.
Acknowledgments
This work was supported in part by the Brazilian Research Council CNPq 305617/2021-4.
Appendix A: The spectrum of $H^{(+)}$ for $L=3$
In table (3) we present the Bethe roots associated with the spectrum of the Hamiltonian $H^{(+)}$ for $L=3$ and for the sectors $Q=0,1$. The energy eigenvalues for sector $Q=2$ are the same as those of sector $Q=1$. Recall that the Bethe roots for $Q=2$ are obtained from the data of the sector $Q=1$ by reflecting the roots around the origin.
$E^{(+)}$ $Q$ $s_{p}$ $\lambda_{j}$
-7.99554373 0 0 $09225911\pm 0.53094089i0.0$$09225911\pm 0.53094089i$
-6.10495278 1 $\frac{1}{3}$ $0.12712369\pm 0.54308217i,~0.04043704\pm 0.53625306i,~0.26682758$
-4.92871242 1 $\frac{2}{3}$ $0.22896922\pm 0.49856802i,~0.011447035\pm 0.51625714i$0.17891715-0.17891715
[…]
-2.57115043 1 $\frac{5}{3}$ $0.15570094\pm 0.53580121i$$23995930\pm 0.50494266i0.0$21352450.02135245
-1.40303732 0 0 $\pm 0.52006713i$$\pm 0.18946913$
-1.36808057 1 $\frac{2}{3}$ $0.06125792\pm 0.52659924i$$0.29627169$, -$0.10619000$, -$0.65151671+i\frac{\pi}{2}$
[…]
0.96180183 0 1 $16995935\pm 0.53302564i0.0$08617700.008617700.199796050.19979605
0.96180183 0 1-1 $16995935\pm 0.53302564i0.0$0861770-0.008617700.19979605-0.19979605
[…]
2.47037783 0 0 $\pm 0.09118002$$\pm 1.07984698i$
3.60850280 1 $\frac{5}{3}$ $0.15793948\pm 0.75208457i$0.233696920.233696920.003285660.003285660.19352258-0.19352258
3.93923101 1 $\frac{1}{3}$ $0.21465852\pm 0.95822988i$0.326699080.326699080.073938240.073938240.09692341-0.09692341
[…]
Table 3: The eigenvalues ($E$) of the Hamiltonian $H^{(+)}$ (10) for $L=3$ together with respective sector ($Q$) and the values of the Bethe roots ($\lambda_{j}$). The Bethe roots and the spin ($s_{p}$) for the sector $Q=2$ have the opposite values of those of sector $Q=1$.
Appendix B: The spectrum of $H^{(c)}$ for $L=3$
In tables (4,5) we present the Bethe roots associated with the spectrum of the Hamiltonian $H^{(c)}$ for $L=3$.
$E^{(c)}$ $\nu_{c}$ $s_{p}$ $\lambda_{j}$
-8.53674848 ++ 0 $0.20734291\pm 0.56486237i$$0.20734291\pm 0.56486237i$$\pm 0.54454434i$
-7.28852460 ++ 0 $08974272\pm 0.51111514i0.0$$08974272\pm 0.51111514i$$\pm 0.58987350$
-5.60503415 - $\frac{1}{2}$ $\begin{array}[c]{c}0.20589826\pm 0.53508028i,~-0.00365705\pm 0.52532788i\\ ~-0.21805593,~-0.18642647+i\frac{\pi}{2}\end{array}$
-5.60503415 - $\frac{1}{2}$ $\begin{array}[c]{c}-0.20589826\pm 0.53508028i,~0.00365705\pm 0.52532788i\\ ~0.21805593,~0.18642647+i\frac{\pi}{2}\end{array}$
-2.99101531 ++ 11 $\begin{array}[c]{c}0.30208439\pm 0.62106348i,~0.06639032\pm 0.53501552i\\ -0.10649764,~-0.63045180\end{array}$
-2.99101531 ++ 1-1 $\begin{array}[c]{c}-0.30208439\pm 0.62106348i,~-0.06639032\pm 0.53501552i\\ 0.10649764,~0.63045180\end{array}$
[…]
-2.44884573 ++ 0 $\pm 0.52250539i$$\pm 1.14014422i$$\pm 0.21314320$
-2.43937374 ++ 11 $\begin{array}[c]{c}0.09827709\pm 0.51654092i,~-0.36360829\pm 0.61547441i\\ 0.60739795,~-0.07673556\end{array}$
-2.43937374 ++ 1-1 $\begin{array}[c]{c}-0.09827709\pm 0.51654092i,~0.36360829\pm 0.61547441i\\ -0.60739795,~0.07673556\end{array}$
-0.91370050 - $\frac{1}{2}$ $\begin{array}[c]{c}0.03789187\pm 0.52617360i,~0.24829909,~-0.12516643\\ ~-0.65421390,~0.45529749+i\frac{\pi}{2}\end{array}$
-0.91370050 - $\frac{1}{2}$ $\begin{array}[c]{c}-0.03789187\pm 0.52617360i,~-0.24829909,~0.12516643\\ ~0.65421390,~-0.45529749+i\frac{\pi}{2}\end{array}$
Table 4: The eigenvalues ($E^{(c)}$) of the Hamiltonian $H^{(c)}$ (14) for $L=3$. For each energy, we indicate the sector ($\nu_{c}$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$).
$E^{(c)}$ $\nu_{c}$ $s_{p}$ $\lambda_{j}$
0 - $\frac{3}{2}$ $\begin{array}[c]{c}-0.11216447\pm 0.51884900i,~0.26708111,~0.05141273\\ ~-0.62612820,~0.53196329+i\frac{\pi}{2}\end{array}$
0 - $\frac{3}{2}$ $\begin{array}[c]{c}0.11216447\pm 0.51884900i,~-0.26708111,~-0.05141273\\ ~0.62612820,~-0.53196329+i\frac{\pi}{2}\end{array}$
0.43937374 ++ 11 $\begin{array}[c]{c}0.09986159\pm 1.07140349i,~-0.20825039\pm 0.52974992i\\ 0.21482995,~0.00194763\end{array}$
0.43937374 ++ 1-1 $\begin{array}[c]{c}-0.09986159\pm 1.07140349i,~0.20825039\pm 0.52974992i\\ -0.21482995,~-0.00194763\end{array}$
2.14093253 - $\frac{1}{2}$ $\begin{array}[c]{c}-0.2309567\pm 0.52156409i,~0.67763416,~0.13923161\\ ~-0.02538219,~-0.32957012+i\frac{\pi}{2}\end{array}$
2.14093253 - $\frac{1}{2}$ $\begin{array}[c]{c}0.2309567\pm 0.52156409i,~-0.67763416,~-0.13923161\\ ~0.02538219,~0.32957012+i\frac{\pi}{2}\end{array}$
[…]
2.85105230 ++ 0 $\pm 0.75371919i$$\pm 0.62598793$$\pm 0.09437378$
[…]
4.37780211 - $\frac{1}{2}$ $\begin{array}[c]{c}0.41477805\pm 0.61080054i,~0.09748499,~-0.06279076\\ ~-0.28298807,~-0.58126226+i\frac{\pi}{2}\end{array}$
4.37780211 - $\frac{1}{2}$ $\begin{array}[c]{c}-0.41477805\pm 0.61080054i,~-0.09748499,~0.06279076\\ ~0.28298807,~0.58126226+i\frac{\pi}{2}\end{array}$
4.99101531 ++ 11 $\begin{array}[c]{c}-0.28591401\pm 1.02997584i,~0.68123228,~0.14138378\\ -0.02346981,~-0.22731814\end{array}$
4.99101531 ++ 1-1 $\begin{array}[c]{c}0.28591401\pm 1.02997584i,~-0.68123228,~-0.14138378\\ 0.02346981,~0.22731814\end{array}$
[…]
6.22159457 ++ 0 $\pm 0.29687041$$\pm 0.07515995$$\pm 0.61636095+i\frac{\pi}{2}$
Table 5: The eigenvalues ($E^{(c)}$) of the Hamiltonian $H^{(c)}$ (14) for $L=3$. For each energy, we indicate the sector ($\nu_{c}$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$).

### v2

@@
 Keywords: three-state Potts quantum chain, twisted boundaries, Bethe equations.
-December  2025
+January  2026
 1 Introduction
@@
 This operator is usually denominated the diagonal-to-diagonal transfer matrix and its matrix elements are given by
-$\left[T_{\mathrm{dia}}(x)\right]^{b_{1},\dots,b_{L}}_{a_{1},\dots,a_{L}}=\prod_{j=1}^{L}W_{v}(a_{j},b_{j}|x)W_{h}(a_{j},b_{j+1}|x),$ (3)
+$\left[T_{dia}(x)\right]^{b_{1},\dots,b_{L}}_{a_{1},\dots,a_{L}}=\prod_{j=1}^{L}W_{v}(a_{j},b_{j}|x)W_{h}(a_{j},b_{j+1}|x),$ (3)
 where $b_{L+1}=b_{1}$ for periodic boundary conditions. For the three-state scalar Potts model, the elements $W_{h,v}(i,j|x)$ are the entries of the matrices defined by Eqs.(1,2).
@@
 We investigate the solutions of the Bethe equations for $L=2$ and we use the Bethe roots to compute the Hamiltonian energies and the momenta of the eigenstates.
+The respective data for the eigenvalues of the transfer matrix operators are also presented.
 In section 5, we use our framework to build integrable Hamiltonians with periodic boundary conditions whose spectra are obtained by combining the eigenvalues of the spin chains with different types of toroidal boundary conditions.
@@
 It is well known that an important ingredient for the formulation of the integrability of a given vertex model is the corresponding Lax operator [20].
+This operator contains the information about the structure of the Boltzmann weights of the vertex model denoted here by $\mathbb{L}_{i,j}^{k,l}(x)$.
+The indices $i,j,k,l$ represent the four state variables assigned to the links of the square lattice meeting together at a given vertex.
 In this sense, it has been argued that such an operator can be constructed from the spin model edge weights as follows [19],
-$\mathbb{L}_{12}(x)=\sum_{i,j,k=1}^{n}W_{h}(j,i|x)W_{v}(j,k|x)e_{ik}\otimes e_{ji},$ (17)
-where $e_{i,j}$ denotes the standard $n\times n$ Weyl matrices.
+$\mathbb{L}_{12}(x)=\sum_{i,j,k,l=1}^{n}\mathbb{L}_{i,j}^{k,l}(x)~e_{ik}\otimes e_{jl}=\sum_{i,j,k,l=1}^{n}W_{h}(j,i|x)W_{v}(j,k|x)\delta_{i,l}~e_{ik}\otimes e_{jl},$ (17)
+where $e_{i,j}$ denotes the standard $n\times n$ Weyl matrices and the Lax operator is schematically illustrated in Fig.(1).
+[…]
+Figure 1: The illustration of the elements of the Lax operator (17) on the square lattice.
 In addition to that, it has been shown that one can exhibit a $\mathrm{R}$-matrix expressed also in terms of the spin edge weights,
@@
 An immediate consequence of this symmetry is the possibility to define a family of commuting transfer matrices as follows,
-$\mathrm{T}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathbb{L}_{{\cal{A}}1}(x)\right],$ (21)
-where the symbol $\cal{A}$ denotes the auxiliary space $\mathbb{C}^{n}$. We note that the periodic boundary condition corresponds to choosing $\mathrm{G}$ as the $n\times n$ identity matrix.
+$T_{ver}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathbb{L}_{{\cal{A}}1}(x)\right],$ (21)
+where the symbol $\cal{A}$ denotes the auxiliary space $\mathbb{C}^{n}$.
+We note that the periodic boundary condition corresponds to choosing $\mathrm{G}$ as the $n\times n$ identity matrix.
+The matrix elements $T_{ver}(x)]_{a_{1}\dots a_{L}}^{b_{1}\dots b_{L}}$ of the transfer matrix are illustrated in Fig.(2) where $G(a,b)$ denotes the entries of the boundary seam $G$.
+$\cdots$$T_{ver}(x)]_{a_{1}\dots a_{L}}^{b_{1}\dots b_{L}}$$\sum_{c_{1},\dots,c_{L+1}=1}^{n}$$a_{1}$$a_{2}$$a_{L-1}$$a_{L}$$b_{1}$$b_{2}$$b_{L-1}$$b_{L}$$c_{1}$$c_{2}$$c_{3}$$c_{L-1}$$c_{L}$$c_{L+1}$$G(c_{L+1},c_{1})$
+Figure 2: Schematic representation of the elements of the row-to-row transfer matrix of vertex models with twisted boundary $G$.
+As usual the partition function $Z_{ver}(L)$ of the vertex model on the $L\times L$ square lattice is given by the trace of the $L$-th power of the row-to-row transfer matrix,
+$Z_{ver}(L)=\mathrm{Tr}_{{\cal V}}\left[\Big(T_{ver}(x)\Big)^{L}\right].$ (22)
+where ${\cal V}=\displaystyle\prod_{j=1}^{L}\otimes{\mathrm{C}}^{n}$ is Hilbert space associated to the underlying quantum spin chain.
 To define the corresponding quantum spin chain, we assume that the edge weights fulfill the following initial conditions,
@@
 $\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}=P_{j,j+1}\frac{\mathrm{d}}{\mathrm{d}x}\mathbb{L}_{j,j+1}(x)\Big|_{x=0},~~~\mathrm{H}^{(\mathrm{bound})}_{L1}=\mathrm{G}_{L}^{-1}\mathrm{H}^{(\mathrm{bulk})}_{L,1}\mathrm{G}_{L}.$ (25)
-At this point, we also note that $\mathrm{T}(0)$ plays the role of a translation operator which shifts the spin generators by one site modulo $L$.
-More precisely, one can verify the following relations,
-$\mathrm{T}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}\right]\mathrm{T}^{-1}(0)=\mathrm{H}^{(\mathrm{bulk})}_{j+1,j+2},~~j=1,\dots,L-2,$
-$\mathrm{T}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{L-1,L}\right]\mathrm{T}^{-1}(0)=\mathrm{H}^{(\mathrm{bound})}_{L,1}$. (25)
+At this point, we also note that $T_{ver}(0)$ plays the role of a translation operator which shifts the spin generators by one site modulo $L$. More precisely, one can verify the following relations,
+$T_{ver}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}\right]\left[T_{ver}(0)\right]^{-1}=\mathrm{H}^{(\mathrm{bulk})}_{j+1,j+2},~~j=1,\dots,L-2,$
+$T_{ver}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{L-1,L}\right]\left[T_{ver}(0)\right]^{-1}=\mathrm{H}^{(\mathrm{bound})}_{L,1}$. (26)
 Let us now discuss the above construction in the specific case of the three-state scalar Potts model.
@@
 and as a result, we can define three families of commuting transfer matrices,
-$\mathrm{T}^{(\pm)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(\pm)}_{{\cal A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right],~~~\mathrm{T}^{(c)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(c)}_{\cal{A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right].$ (27)
-The transfer matrices $\mathrm{T}^{(\pm)}(x)$ are invariant under the $Z(3)$ spin rotation and their Hamiltonian limits, apart from a constant bulk term, correspond to the quantum spin chains $H^{(\pm)}$.
+$T_{ver}^{(\pm)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(\pm)}_{{\cal A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right],~~T_{ver}^{(c)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(c)}_{\cal{A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right].$ (28)
+The transfer matrices $T_{ver}^{(\pm)}(x)$ are invariant under the $Z(3)$ spin rotation and their Hamiltonian limits, apart from a constant bulk term, correspond to the quantum spin chains $H^{(\pm)}$.
 Thus, we have the following commutation relations,
-$\mathrm{T}^{(\pm)}(x),{\cal{O}}_{Z(3)}]=0,~~[\mathrm{T}^{(\pm)}(x),H^{(\pm)}]=0$ (28)
-such that the spectra of $\mathrm{T}^{(+)}(x)$ and $\mathrm{T}^{(-)}(x)$ are related by the following equivalences,
-$\mathrm{Spec}\big[T^{(+)}\big]_{Q=0}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=0},~~\mathrm{Spec}\big[T^{(+)}\big]_{Q=1}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=2},~~\mathrm{Spec}\big[T^{(+)}\big]_{Q=2}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=1}$ (29)
-The transfer matrix $\mathrm{T}^{(c)}(x)$ keeps only the invariant concerning the $Z(2)$ subgroup associated with the charge conjugation.
+$T_{ver}^{(\pm)}(x),{\cal{O}}_{Z(3)}]=0,~~[T_{ver}^{(\pm)}(x),H^{(\pm)}]=0$ (29)
+such that the spectra of $T_{ver}^{(+)}(x)$ and $T_{ver}^{(-)}(x)$ are related by the following equivalences,
+$\mathrm{Spec}\big[T_{ver}^{(+)}\big]_{Q=0}=\mathrm{Spec}\big[T_{ver}^{(-)}\big]_{Q=0},~~\mathrm{Spec}\big[T_{ver}^{(+)}\big]_{Q=1}=\mathrm{Spec}\big[T_{ver}^{(-)}\big]_{Q=2},~~\mathrm{Spec}\big[T_{ver}^{(+)}\big]_{Q=2}=\mathrm{Spec}\big[T_{ver}^{(-)}\big]_{Q=1}.$ (30)
+The transfer matrix $T_{ver}^{(c)}(x)$ keeps only the invariant concerning the $Z(2)$ subgroup associated with the charge conjugation.
 The respective Hamiltonian limit gives rise to the quantum spin chain $H^{(c)}$, namely
-$\mathrm{T}^{(c)}(x),{\cal{O}}_{Z(2)}]=0,~~[\mathrm{T}^{(c)}(x),H^{(c)}]=0$ (30)
+$T_{ver}^{(c)}(x),{\cal{O}}_{Z(2)}]=0,~~[T_{ver}^{(c)}(x),H^{(c)}]=0$ (31)
 In the next sections, we shall see that the construction of these transfer matrices is essential to set up the Bethe equations for the corresponding quantum spin chains.
-From the spectral relations (29) it is enough to consider the problem for $T^{(+)}(x)$ and $T^{(c)}(x)$.
-3 The Bethe equation for $T^{(+)}(x)$
+From the spectral relations (30) it is enough to consider the problem for $T_{ver}^{(+)}(x)$ and $T_{ver}^{(c)}(x)$.
+3 The Bethe equation for $T_{ver}^{(+)}(x)$
 Here, we follow the strategy devised by Baxter to solve the eigenvalue problem associated with the diagonal-to-diagonal transfer matrix of the Ising model [1].
 We recall that this kind of approach has also been used to uncover the Bethe equation for the critical $Z(n)$ spin model with periodic boundary conditions [11, 22].
-The first step is to build a suitable proposal for the form of the eigenvalues of the transfer matrix $T^{(+)}$.
+The first step is to build a suitable proposal for the form of the eigenvalues of the transfer matrix $T_{ver}^{(+)}$.
 To this end, we observe any matrix elements of such operator have the following structure,
-$\left[T^{(+)}\right]_{a_{1},\dots,a_{L}}^{b_{1},\dots,b_{L}}=\Big[\frac{1}{g(x)g_{1}(x)}\Big]^{L}H(x)$ (31)
+$\left[T_{ver}^{(+)}\right]_{a_{1},\dots,a_{L}}^{b_{1},\dots,b_{L}}=\Big[\frac{1}{g(x)g_{1}(x)}\Big]^{L}H(x)$ (32)
 where $g(x)$ and $g_{1}(x)$ take into account the denominators of the edge weights,
@@
 Considering that the transfer matrix commutes for distinct values of the spectral parameter, we can write the eigenvalue problem as follows,
-$T^{(+)}(x)\left|\phi\right\rangle=\Lambda^{(+)}(x)\left|\phi\right\rangle$ (34)
+$T_{ver}^{(+)}(x)\left|\phi\right\rangle=\Lambda^{(+)}(x)\left|\phi\right\rangle$ (35)
 where the eigenstates $\left|\phi\right\rangle$ do not depend on $x$.
-Due to the independence of the eigenstates on $x$ a given eigenvalue $\Lambda^{(+)}(x)$ can be regarded as a linear combination of the elements of the transfer matrix $T^{(+)}(x)$ with constant coefficients.
-As a consequence, we can conclude that $\Lambda^{(+)}(x)$ should have the same form as given by Eqs.(31,33).
+Due to the independence of the eigenstates on $x$ a given eigenvalue $\Lambda^{(+)}(x)$ can be regarded as a linear combination of the elements of the transfer matrix $T_{ver}^{(+)}(x)$ with constant coefficients.
+As a consequence, we can conclude that $\Lambda^{(+)}(x)$ should have the same form as given by Eqs.(32,34).
 To obtain further insights on the structure of $\Lambda^{(+)}(x)$ we have studied the eigenvalue problem (35) for small finite sizes.
@@
 Here we have been inspired by the matrix structure derived for the diagonal-to-diagonal transfer matrix of the three-state Chiral Potts model [12, 13].
-We have found that the transfer matrix $T^{(+)}(x)$ satisfies the following identity,
-$T^{(+)}(x-\frac{\pi}{3})T^{(+)}(x-\frac{\pi}{6})T^{(+)}(x)$ $=$ $T^{(+)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T^{(+)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T^{(+)}(x)$ (37)
-$+$ $\big[f_{3}(x)\big]^{L}T^{(+)}(x+\frac{\pi}{3})\bigg\}$
+We have found that the transfer matrix $T_{ver}^{(+)}(x)$ satisfies the following identity,
+$T_{ver}^{(+)}(x-\frac{\pi}{3})T_{ver}^{(+)}(x-\frac{\pi}{6})T_{ver}^{(+)}(x)$ $=$ $T_{ver}^{(+)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T_{ver}^{(+)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T_{ver}^{(+)}(x)$ (38)
+$+$ $\big[f_{3}(x)\big]^{L}T_{ver}^{(+)}(x+\frac{\pi}{3})\bigg\}$
 where the auxiliary functions $f_{1}(x)$, $f_{2}(x)$ and $f_{3}(x)$ are determined by,
 $f_{1}(x)=3\tan(x)\cot(x+\frac{\pi}{6}),~~f_{2}(x)=3\tan(x-\frac{\pi}{6})\cot(x),~~f_{3}(x)=3\tan(x-\frac{\pi}{6})\cot(x+\frac{\pi}{6})$ (39)
+The way we have established the functional relation (38,39) is as follows.
+We first solve analytically some suitable entries of this matrix relation for $L=2$ and $L=3$ allowing the determination of the coefficients $f_{1}(x)$, $f_{2}(x)$ and $f_{3}(x)$.
+With the help of symbolic algebra packages we have verified that all the entries of the functional relation are indeed satisfied up to $L=5$.
+In this task we find convenient to transform the trigonometric functions in terms of ratio of polynomials in the variable $\exp(2ix)$.
+Based on this checking we conjecture that Eqs.(38,39) should be valid for arbitrary $L$ but a systematic proof of that has eluded us so far.
 We can turn the above matrix equation into a functional relation for the eigenvalues by acting on both sides of Eq.(38) on a common set of eigenvectors.
@@
 $E_{Q}^{(+)}=-\frac{1}{\Lambda^{(+)}(0)}\frac{\mathrm{d}}{\mathrm{d}x}\Lambda^{(+)}(x)\Big|_{x=0}-4L\cot(\frac{\pi}{3})=\sum_{j=1}^{\bar{N}_{Q}}\cot\Big(\frac{\pi}{6}-\xi_{j}\Big)+i\mu_{Q}-\frac{2L}{\sqrt{3}}$ (42)
-At this point, we note that Eqs.(40,41) leads us to the expressions of the Bethe equation (11) and the energy eigenvalue (12) presented in the introduction by choosing,
+At this point, we note that Eqs.(41,42) leads us to the expressions of the Bethe equation (11) and the energy eigenvalue (12) presented in the introduction by choosing,
 $\xi_{j}=i\lambda_{j}+\frac{\pi}{12}$ (43)
@@
 We now describe the set of roots solving the Bethe equation (11) associated with the energy eigenvalues of the Hamiltonian $H^{(+)}$ for $L=2$.
-We obtain some insight on the nature of the roots, exploring the fact that the eigenvectors of $T^{(+)}(x)$ are independent of the spectral parameter.
-By choosing a suitable value for $x$ we can compute numerically the eigenvectors making it possible the determination of the corresponding eigenvalues as polynomial functions on $x$.
+We obtain some insight on the nature of the roots, exploring the fact that the eigenvectors of $T_{ver}^{(+)}(x)$ are independent of the spectral parameter.
+By choosing a suitable value for $x$ we can compute numerically the eigenvectors making it possible the determination of the corresponding eigenvalues as polynomial functions on the variable $\exp(2ix)$.
 The zeros of such polynomials work as starting values for the numerical solution of the Bethe equation (11) providing us the Bethe roots with high precision.
@@
 The energy values are obtained with the help of Eqs.(12,13) and we have verified that they agree with those obtained from the exact diagonalization of the Hamiltonian $H^{(+)}$.
+We remark that these Bethe roots allows us to compute the corresponding transfer matrix eigenvalues with the the help of Eqs.(37,43) for arbitrary values of the spectral parameter $x$.
+As an example we have added in the last column of table (1) the eigenvalues $\Lambda^{(+)}(x)$ at the point $x=\frac{3}{5}$ reproducing the values attained from the diagonalization of the operator $T_{ver}^{(+)}(\frac{3}{5})$.
 From table (1) we note that if a given energy in sector $Q=1$ is parametrized by a set of roots $\{\lambda_{j}\}$ the same energy is obtained in sector $Q=2$ using the opposite set of roots $\{-\lambda_{j}\}$.
-From our solution, we see that this property is valid for arbitrary $L$ in accordance with the fact that the spectra for sectors $Q=1,2$ are the same.
-In addition, from Eqs.(13,43) we see that the momenta of such states are however opposite to each other.
-$E^{(+)}$ $Q$ $s_{p}$ $\lambda_{j}$
--4.93624921 0 0 $\pm 0.53202156i$
--2.30940107 1 $\frac{1}{3}$ $0.06315064\pm 0.54254935i,~0.19822180$
--2.30940107 2 $\frac{1}{3}$ $0.06315064\pm 0.54254935i$0.19822180-0.19822180
--0.84529946 1 $\frac{2}{3}$ $0.16532518\pm 0.50188873i$0.10277717-0.10277717
--0.84529946 2 $\frac{2}{3}$ $0.16532518\pm 0.50188873i,~0.10277717$
-1.15470053 0 1 $0,~\frac{i\pi}{2}$
-3.15470053 1 $\frac{2}{3}$ 0.245380170.245380170.02638725-0.026387250.64959867-0.64959867
-3.15470053 2 $\frac{2}{3}$ 0.24538017-0.245380170.02638725,0.649598670.02638725,~0.64959867
-3.78154867 0 0 $\pm 0.12258177$
+From our solution, we see that this property is valid for arbitrary $L$ in accordance with the fact that the Hamiltonian spectra for sectors $Q=1,2$ are the same since besides the spectral relation (30) we have that the invariance under reflection maps $H^{(+)}$ to $H^{(-)}$.
+We note that such spectral degeneracy does not apply to the transfer matrix eigenvalues due to the fact that we only have the spectral equivalence (30) for generic values of the spectral parameter.
+In addition, from Eqs.(13,44) we see that the momenta of the states with roots $\{\lambda_{j}\}$ and $\{-\lambda_{j}\}$ are however opposite to each other.
+$E^{(+)}$ $Q$ $s_{p}$ $\lambda_{j}$ $\Lambda^{(+)}(\frac{3}{5})$
+-4.93624921 0 0 $\pm 0.53202156i$ $0.95403776$
+-2.30940107 1 $\frac{1}{3}$ $0.06315064\pm 0.54254935i,~0.19822180$ $\begin{array}[c]{c}1.16822213\\ +0.04919793i\end{array}$
+-2.30940107 2 $\frac{1}{3}$ $0.06315064\pm 0.54254935i$0.19822180-0.19822180 $\begin{array}[c]{c}1.16822213\\ -0.04919793i\end{array}$
+-0.84529946 1 $\frac{2}{3}$ $0.16532518\pm 0.50188873i$0.10277717-0.10277717 $\begin{array}[c]{c}1.32149657\\ -0.07557454i\end{array}$
+-0.84529946 2 $\frac{2}{3}$ $0.16532518\pm 0.50188873i,~0.10277717$ $\begin{array}[c]{c}1.32149657\\ +0.07557454i\end{array}$
+1.15470053 0 1 $0,~\frac{i\pi}{2}$ $1.54404018$
+3.15470053 1 $\frac{2}{3}$ 0.245380170.245380170.02638725-0.026387250.64959867-0.64959867 $\begin{array}[c]{c}1.79532310\\ +0.025796612i\end{array}$
+3.15470053 2 $\frac{2}{3}$ 0.24538017-0.245380170.02638725,0.649598670.02638725,~0.64959867 $\begin{array}[c]{c}1.79532310\\ -0.025796612i\end{array}$
+3.78154867 0 0 $\pm 0.12258177$ $1.88075838$
 Table 1: The eigenvalues ($E^{(+)}$) of the Hamiltonian $H^{(+)}$ (10) for $L=2$.
 For each energy, we indicate the sector ($Q$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$).
+We have also added the respective transfer matrix eigenvalue $\Lambda^{(+)}(x)$ at $x=\frac{3}{5}$.
 Inspecting table (1) we see the occurrence of low-lying eigenstates whose values for the spin are $\pm\frac{1}{3}$ and $\pm\frac{2}{3}$.
@@
 $\bar{N}_{0}=2L-2,~~\bar{N}_{1}=\bar{N}_{2}=2L-1$ (47)
-4 The Bethe equation for $T^{(c)}(x)$
+4 The Bethe equation for $T_{ver}^{(c)}(x)$
 Here, we use the same procedure explained in the previous section to determine the parameterization of the spectrum of Hamiltonian $H^{(c)}$ in terms of a Bethe ansatz equation.
 We therefore avoid the repetition of technical details and concentrate our presentation to the main results.
-In particular, we find out that the structure of the eigenvalues of $T^{(c)}(x)$ can be given as follows,
+In particular, we find out that the structure of the eigenvalues of $T_{ver}^{(c)}(x)$ can be given as follows,
 $\Lambda^{(c)}(x)=\Big[\frac{g(\frac{\pi}{6})g_{1}(\frac{\pi}{6})}{g(x)g_{1}(x)}\Big]^{L}\prod_{k=1}^{2L}\frac{\sin(\xi_{k}-\frac{\pi}{6}+x)}{\sin(\xi_{k})}$ (48)
 where now the number of sine terms is the same as that present in the general form (34) for the transfer matrix elements.
-We arrive at this conclusion by inspecting the asymptotic limit of the eigenvalues of $T^{(c)}(x)$ at the points $\pm i\infty$ for small lattice sizes.
+We arrive at this conclusion by inspecting the asymptotic limit of the eigenvalues of $T_{ver}^{(c)}(x)$ at the points $\pm i\infty$ for small lattice sizes.
 As before, the Bethe equation for the variables $\xi_{k}$ is uncovered with the help of a functional relation for the eigenvalue involving distinct spectral parameters.
 This equation is derived from a matrix identity for the product of transfer matrices whose main structure is similar to that presented in previous section, see Eq.(38).
-More specifically, we find that $T^{(c)}(x)$ satisfies the following matrix identity,
-$T^{(c)}(x-\frac{\pi}{3})T^{(c)}(x-\frac{\pi}{6})T^{(c)}(x)$ $=$ $T^{(c)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T^{(c)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T^{(c)}(x)$ (48)
-- $\big[f_{3}(x)\big]^{L}T^{(c)}(x+\frac{\pi}{3})\bigg\}$
+More specifically, we find that $T_{ver}^{(c)}(x)$ satisfies the following matrix identity,
+$T_{ver}^{(c)}(x-\frac{\pi}{3})T_{ver}^{(c)}(x-\frac{\pi}{6})T_{ver}^{(c)}(x)$ $=$ $T_{ver}^{(c)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T_{ver}^{(c)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T_{ver}^{(c)}(x)$ (49)
+- $\big[f_{3}(x)\big]^{L}T{ver}^{(c)}(x+\frac{\pi}{3})\bigg\}$
 where the main difference with Eq.(38) is the presence of a minus sign on the term proportional to the amplitude $f_{3}(x)$.
+We have established this relation using the same procedure explained in the previous section.
+We have checked that such matrix relation is satisfied up to $L=5$ and we conjecture that it is valid for arbitrary $L$.
 If we choose $x=\frac{\pi}{6}-\xi_{k}$ the above matrix identity can turn into a constraint for the eigenvalues (48) producing a set of relations for the variables $\xi_{k}$.
@@
 If we denote these eigenvalues by the symbol $\nu_{c}$ we find that there are $(3^{L}+1)/2$ states with $\nu_{c}=+1$ and $(3^{L}-1)/2$ states with $\nu_{c}=-1$.
-We obtain the initial values for the Bethe roots by computing the eigenvalues of the transfer matrix $T^{(c)}(x)$ as meromorphic functions on $x$.
-We use this data to solve the Bethe equation (15) with high numerical precision and the energy and the spin of the states are calculated with the help of Eqs.(16,51).
+We obtain the initial values for the Bethe roots by computing the eigenvalues of the transfer matrix $T_{ver}^{(c)}(x)$ as meromorphic functions on $x$.
+We use this data to solve the Bethe equation (15) with high numerical precision and the energy and the spin of the states are calculated with the help of Eqs.(16,52).
 The energies obtained from the Bethe roots are then checked with those coming from the exact diagonalization of the Hamiltonian $H^{(c)}$.
 The results are shown in table (2).
-$E^{(c)}$ $\nu_{c}$ $p$ $\lambda_{j}$
--5.77350269 +1 0 $0.13588376\pm 0.57005210i$$0.13588376\pm 0.57005210i$
--3.89397505 +1 0 $\pm 0.50828016i,~\pm 0.524244017$
-[…]
-[…]
-1.154700538 +1 11 $0.28844398\pm 0.61762004i$0.558766300.558766300.018121660.01812166
-1.154700538 +1 1-1 $0.28844398\pm 0.61762004i$0.55876630-0.558766300.01812166-0.01812166
-2.739274521 +1 0 $\pm 1.08877871i$$\pm 0.14111440$
-[…]
-[…]
+In this table we have also presented the data for the corresponding transfer matrix eigenvalue $\Lambda^{(c)}(x)$ at $x=\frac{3}{5}$ computed by using Eqs.(43,48).
+We stress that these transfer matrix eigenvalues agree with the values obtained from the diagonalization of the operator $T_{ver}^{(c)}(\frac{3}{5})$.
+$E^{(c)}$ $\nu_{c}$ $p$ $\lambda_{j}$ $\Lambda^{(c)}(\frac{3}{5})$
+-5.77350269 +1 0 $\begin{array}[c]{c}0.13588376\pm 0.57005210i,\\ -0.13588376\pm 0.57005210i\end{array}$ $0.90168671$
+-3.89397505 +1 0 $\pm 0.50828016i,~\pm 0.524244017$ $1.02818296$
+-1.67372658 1-1 $\frac{1}{2}$ $\begin{array}[c]{c}0.13542121\pm 0.5300866i,\\ -0.14584486,-0.12499756+i\frac{\pi}{2}\end{array}$ $\begin{array}[c]{c}1.23423027\\ -0.06662064i\end{array}$
+-1.67372658 1-1 $\frac{1}{2}$ $\begin{array}[c]{c}-0.13542121\pm 0.5300866i,\\ 0.14584486,~0.12499756+i\frac{\pi}{2}\end{array}$ $\begin{array}[c]{c}1.23423027\\ +0.06662064i\end{array}$
+1.154700538 +1 11 $\begin{array}[c]{c}-0.28844398\pm 0.61762004i,\\ 0.55876630,~0.01812166\end{array}$ $\begin{array}[c]{c}1.54404018\\ +0.03665078i\end{array}$
+1.154700538 +1 1-1 $\begin{array}[c]{c}0.28844398\pm 0.61762004i,\\ -0.55876630,~-0.01812166\end{array}$ $\begin{array}[c]{c}1.54404018\\ -0.03665078i\end{array}$
+2.739274521 +1 0 $\pm 1.08877871i$$\pm 0.14111440$ $1.73331161$
+3.983127663 1-1 $\frac{1}{2}$ $\begin{array}[c]{c}-0.59441775,~-0.05108087,\\ ~0.19398856,~0.45151005+i\frac{\pi}{2}\end{array}$ $\begin{array}[c]{c}1.91132869\\ +0.00914203i\end{array}$
+3.983127663 1-1 $\frac{1}{2}$ $\begin{array}[c]{c}0.59441775$,~$0.05108087,\\ -0.19398856,~-0.45151005+i\frac{\pi}{2}\end{array}$ $\begin{array}[c]{c}1.91132869\\ -0.00914203i\end{array}$
 Table 2: The eigenvalues ($E^{(c)}$) of the Hamiltonian $H^{(c)}$ (14) for $L=2$.
 For each energy, we indicate the sector ($\nu_{c}$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$).
+We have also added the respective transfer matrix eigenvalue $\Lambda^{(c)}(x)$ at $x=\frac{3}{5}$.
 Interestingly enough, we observe that there are excitations with spin $\pm\frac{1}{2}$ in the charge sector with negative eigenvalue.
@@
 We can explore the invariance of the Yang-Baxter algebra (20) to construct more general families of commuting transfer matrices. For example, we can write,
-$\mathrm{\tilde{T}}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}1}(x)\right],$ (53)
+$\tilde{T}_{ver}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}1}(x)\right],$ (54)
 where the boundary seams are spread throughout the bulk by acting them on each of the Lax operators.